An impatient driver, wanting to save time, drives along a city bypass at an average speed of 115 km/h where the speed limit is 90.0 km/h. If the bypass is 18.0 km long, how much time (in minutes) does the driver save by breaking the speed limit? (By the way, the driver consumes about 20% more fuel at the higher speed.)


1
Expert's answer
2021-09-09T10:51:30-0400

The time it will take at 90 and 115 km/h:

"t_1=\frac{18}{90}=0.2\text{ h},\\\space\\
t_1=\frac{18}{115}=0.16\text{ h},\\\space\\
\Delta t=0.2-0.16=0.04\text{ h, or 2.4 min}."